### v1

# Look the Other Way: Designing ‘Positive’ Molecules with Negative Data via Task Arithmetic

Abstract
The scarcity of molecules with desirable properties (i.e., ‘positive’ molecules) is an inherent bottleneck for generative molecule design. To sidestep such obstacle, here we propose molecular task arithmetic: training a model on diverse and abundant negative examples to learn ‘property directions’ – without accessing any positively labeled data – and moving models in the opposite property directions to generate positive molecules. When analyzed on 20 zero-shot design experiments, molecular task arithmetic generated more diverse and successful designs than models trained on positive molecules. Moreover, we employed molecular task arithmetic in dual-objective and few-shot design tasks. We find that molecular task arithmetic can consistently increase the diversity of designs while maintaining desirable design properties. With its simplicity, data efficiency, and performance, molecular task arithmetic bears the potential to become the de-facto transfer learning strategy for de novo molecule design.
1 Introduction
Discovering one drug molecule can take over a decade and billions of dollars1, 2. The first obstacle is charting the ‘chemical space’3 effectively. Chemical space is estimated to contain approximately 1060 drug-like molecules, with scarcity of molecules possessing desirable properties, e.g., bioactivity towards a pharmaceutically relevant target. Generative deep learning has emerged as a revolutionizing technology for drug discovery in the last decade – with the potential to shorten de novo design pipelines from years to weeks 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16.
Transfer learning has been fundamental for generative deep learning in molecule design17. The generative models are first pretrained on large, unlabeled molecule datasets to learn unconditional molecule generation. Next, the pretrained models are fine-tuned using molecules of a target property (high solubility, bioactivity, synthetic accessibility, etc.) and sampled to design compounds satisfying the design task18, 19.
The fine-tuning stage faces an inherent challenge: molecules with desirable properties are not only scarce but also lack structural diversity20. In the case of early stage drug discovery, positive molecules (or ‘hits’) can be as rare as 1% in large, diverse molecule libraries. These molecules are often optimized with small structural edits to explore structure-activity relationships, resulting datasets typically containing a few hundred molecules with low structural diversity21, 22, 23, 24. Here, we target such key data limitations in molecule design by proposing a new strategy for generative deep learning – molecular task arithmetic.
Molecular task arithmetic targets the limited availability of desirable and structurally diverse molecules in drug discovery. By identifying directions in the model weight space via training on abundant and diverse negative molecules, molecular task arithmetic moves pretrained model weights in the desirable property direction to create positive molecules. Editing models with task directions is known as task arithmetic25, and has already powered successful applications in computer vision and natural language processing, e.g., for model merging, multi-objective optimization, and privacy-preserving machine learning25, 26. Here, we explore the untapped potential of task arithmetic as a new transfer learning strategy for molecule design. We first show how molecular task arithmetic unlocks zero-shot molecule design, i.e., generating positive molecules without any positive labeled molecules. Next, we use molecular task arithmetic for dual-objective tasks, and as a drop-in replacement for the traditional supervised finetuning in few-shot applications. Across 20 studied design tasks, molecular task arithmetic surpasses supervised finetuning by creating more diverse and accurate molecules. Molecular task arithmetic is simple, applicable to all models optimized via gradient descent, and bears the potential to be the standard transfer learning pipeline for molecule design. Our key contributions are as follows:
  • We propose a simple and widely-applicable transfer learning strategy – molecular task arithmetic. Molecular task arithmetic tackles the inherent data limitations in de novo design by harnessing negative molecules, a direction explored for the first time here. Across experiments, models trained with molecular task arithmetic design more accurate and diverse molecules.
  • We unlock zero-shot ligand-based de novo design with transfer learning. Since molecular task arithmetic uses only negative molecules to learn task directions, it can be used even when no positively labeled data is available. Models trained with molecular task arithmetic can design molecules fulfilling one or multiple target properties, with minimal side effects on non-targeted properties.
  • We improve few-shot molecule design with traditional transfer learning pipelines. When used to augment the traditional finetuning based on positive molecules, molecular task arithmetic increases the number of diverse and desirable designs across 20 experiments. This shows the potential of molecular task arithmetic to replace the dominant transfer learning practices, by allowng to leverage both positive and negative molecules simultaneously.
Fig. 1: Molecular task arithmetic. (a) SMILES string creation from a molecular graph. Atoms are annotated by the periodic table symbols and double bonds are denoted with ‘=’ sign. Lowercase symbols describe aromatic bonds and brackets annotate branches. (b) Chemical language modeling. A next token prediction task can be defined on SMILES strings using a sequence model such as an LSTM. The learned language model can be sampled for molecule design. (c) Zero-shot design. Molecular task arithmetic learns a task direction in the model weight space by applying finetuning on negative molecules (dashed purple arrow). The learned direction is traversed in the opposite direction (solid orange arrow) for zero-shot molecule generation. (d) Dual-objective design. By learning multiple task directions independently, molecular task arithmetic can train multi-objective models. (e) Few-shot design. Finetuning the zero-shot model with known positive molecules (dashed orange arrow) enables molecular task arithmetic to use all available data for the task.
2 Background
Task arithmetic
In the weight space of a trained model, certain directions correspond to improvements in performance on specific tasks 25. These directions are known as “task vectors”25, 27, 28, 29. In a transfer learning setting, where data for a task $t$ are available, a task vector $\tau^{t}$ can be computed by subtracting the pretrained model weights from the finetuned model weights25:
$\tau^{t}=\theta^{t}_{\text{ft}}-\theta_{\text{pt}},$ (1)
where $\theta^{t}_{\text{ft}}$ and $\theta_{\text{pt}}$ are the flattened weights of the finetuned and of the pretrained model, respectively. The task vector can then be used to modify the behavior of the pretrained model, e.g., subtracting the task vector from the pretrained model weights can yield a model that is worse at the finetuning task $t$:
$\theta_{\text{new}}=\theta_{\text{pt}}-\lambda\tau^{t},$ (2)
where $\lambda$ is a scaling factor that determines the step size in the direction. Termed as “forgetting via negation”25, this approach can remove undesired behavior from deep models, such as toxic language generation25. Task vectors can also be added together to create models with multi-task capabilities (“learning via addition”25):
$\theta_{\text{mobj}}=\theta_{\text{pt}}+\lambda(\tau^{t_{1}}+\tau^{t_{2}}+\tau^{t_{3}}+\dots+\tau^{t_{n}}),$ (3)
where $\tau^{t_{i}}$ is the task vector learned for the task $t^{i}$. Both in Eq. (2) and in Eq. (3), $\lambda$ controls how far the final model lands from the pretrained one in the weight space. While too large $\lambda$ values can have detrimental side effects on model performance, too low values fail to sufficiently edit the model behavior.
Molecular string generation for de novo design
Molecular strings are among the most popular molecular representations for de novo molecule design 30, 31, 32, 33. The most well-established representation is the Simplified Molecular Input Line Entry System (SMILES 31) strings (Fig. 1a). SMILES strings are created by traversing a molecular graph in any chosen direction, starting from any non-hydrogen atom, and annotating atomic symbols, plus branches and bond types. Multiple SMILES strings (often referred to as randomized SMILES) can be created by starting from different atoms (and/or with different traversals); this characteristic allows for data augmentation, where multiple instances referring to the same molecule are used to inflate the amount of training data34. Canonicalization algorithms have been proposed to obtain a single SMILES string from any molecule (i.e., by standardizing the starting atom and the graph traversal rules 35, 36). Canonicalization can aid in duplicate molecule detection and molecule ranking 37, 38, 39. Both versions are commonly used for molecule design, with different advantages and drawbacks 40.
For modeling purposes, the SMILES string represents a molecule $m$ as $\{s_{1},s_{2},\dots,s_{n}\}$, where $s_{i}$ corresponds to the $i^{th}$ token in the molecular string. The next token prediction language modeling objective is then defined as: $\max\sum_{m}^{M}\sum_{i}\log P(s_{i}|s_{<i};\theta)$, where $\theta$ is typically a long-short term memory network41, 42, 14 or a transformer43, 44, 15 (Fig. 1b), and $M$ is a dataset of SMILES strings.
In a transfer learning setting, two phases typically occur: (a) pretraining, where $M$ is selected to contain hundreds of thousands of molecules to learn key elements, such as generating ‘chemically valid’ SMILES, and to learn basic molecular properties17; and (b) (self-)supervised finetuning, where a curated set of molecules with desired properties is used, with the same training objective to condition the model45, 46, 17. New molecules can then be designed by sampling the learned multinomial distribution autoregressively. Chemical language models have already demonstrated powerful design capabilities for drug discovery and are a well-established paradigm for molecule design 47, 48, 4, 49, 40, 14, 15, 16.
3 Molecular Task Arithmetic
Why task arithmetic? Chemical space is vast but sparse: it is estimated to contain 1060 molecules, but molecules of desirable properties are rare 3. Current training strategies rely on these rare, ‘positive’ molecules (i.e., molecules possessing one or more desirable properties), and therefore, the limited data availability and diversity form a bottleneck. Here we ‘look the other way’ and propose a new transfer learning strategy that can leverage the abundant and diverse ‘negative’ molecules, that is, molecular task arithmetic.
Molecular task arithmetic views “forgetting via subtraction” (Eq. 2) as an opportunity to learn only from negative data to design positive molecules. For the task $t$ of designing molecules with a desirable property, molecular task arithmetic finetunes a pretrained model ($\theta_{pt}$) with molecules having non-desirable values of this property. This allows obtaining $\theta_{opp}$, a model that can generate negative molecules. Then the task vector $\tau^{t}$ is computed by subtracting $\theta_{pt}$ from $\theta_{opp}$. Finally, a model that creates the molecules of positive property, $\theta_{task}$ (Fig. 1c), can be constructed by subtracting $\tau^{t}$ from $\theta_{pt}$ with a scaling factor $\lambda$:
$\tau^{t}$ $=\theta_{\text{opp}}-\theta_{\text{pt}},$
$\theta_{\text{task}}$ $=\theta_{\text{pt}}-\lambda\tau^{t}.$ (4)
In other words, molecular task arithmetic learns the ‘direction’ of a molecular property using negative molecules and then ‘looks the other way’ – by moving the pretrained model in the opposite direction to design positive molecules. By not using any labeled positive molecules, this formulation opens the doors for unexplored tasks, e.g., zero-shot molecule design.
To test the impact of task arithmetic on steering de novo design of bespoke molecules, we have applied it to easily computable properties that are relevant for drug discovery, namely: (a) fraction of sp3-hybridized carbons (frac. sp3C), which reflects molecular complexity and 3D characteristics; (b) octanol-water partitioning coefficient (logP), influencing membrane permeability and solubility; (c) number of hydrogen bond donors, which affects a molecule’s ability to form key interactions with biological targets; (d) number of rings (No. Rings), linked to molecular rigidity and target binding specificity; and (e) topological surface area (TPSA).
We test molecular task arithmetic on chemical language models with an LSTM backbone. We pretrained the LSTM on 1.5M drug-like molecules and define conditional molecule design experiments as designing molecules with a higher or lower value than a predefined threshold of the five selected molecular properties (Fig. A1). By defining tasks for randomized and canonical SMILES and per property, we create 20 design tasks. Five splits of 1024 train, 256 validation, 256 test molecules, and 10240 negative molecules are curated per task for finetuning and molecular task arithmetic (see Appendix for more details).
4 Experiments
4.1 Can Task Arithmetic Even Generate Valid Molecules?
The language of SMILES strings possesses a strict grammar to represent molecules, where single-character changes might yield ‘chemically invalid’ strings (i.e., strings that cannot be converted back into molecules). Hence, our first question was whether the capacity to generate chemically valid strings could be retained under the model weight manipulation introduced by task arithmetic.
To answer this question, we applied molecular task arithmetic with $\lambda=0.50$ (Eq. 3) across the ten single-objective design tasks and the two SMILES formats (canonical and randomized). We then computed molecular validity (defined as the ratio of unique and chemically valid design strings among 100,000 generations), as a well-established control of the capacity of a model to learn the ‘chemical language’ of SMILES. Molecular task arithmetic obtains an average validity of 79.0% with randomized SMILES strings and 59.2% with canonical SMILES strings. While these values drop from the pretrained models (which were optimized for validity, obtaining values of 95.2% and 95.8% for randomized and canonical SMILES strings, respectively), molecular task arithmetic preserves its ability to generate chemically meaningful strings, especially with randomized SMILES strings. Remarkably, this finding indicates that the weight space of a chemical language model can be traversed with no loss guidance – opening new and unexplored avenues for de novo molecule design with generative deep learning.
In what follows, we explore the benefits and limitations of traversing the model weight space with molecular task arithmetic, with the following objectives:
  1. 1.
    Exploring zero-shot de novo design, where only molecules with ‘undesired’ properties are used to design ‘desirable’ molecules. In this context, we explore single-property molecule design (Fig. 1c), and dual-objective molecule design (Fig. 1d), where single and pairs of properties are targeted, respectively.
  2. 2.
    Understanding the effect of task vector scaling, where we analyze how the magnitude of the molecular task vector ($\lambda$) affects the capacity to explore the chemical space in search for molecules with desirable properties.
  3. 3.
    Boosting few-shot molecule design (Fig. 1e), by augmenting traditional finetuning (which relies on positive molecules only) with negative molecules, via molecular task arithmetic.
4.2 Zero-Shot Molecule Design
Single-objective design.
Fig. 2: Zero-shot molecule design with molecular task arithmetic. Models were trained on 10 design tasks of randomized SMILES strings, with molecular task arithmetic and supervised finetuning. 100,000 molecules were designed and clustered. (a) The number of cluster centers that possess the desired property, (b) the ratio of the designs that satisfy the design task, and (c) number of clusters were computed. The pretrained model is also included in the analysis as a baseline. Bar heights report the mean statistics across five training sets, and error bars denote the standard deviation. Finetuning was conducted up to the training sets of 1024 molecules, while molecular task arithmetic used 10,240 negative molecules and no labeled positive data.
We benchmarked molecular task arithmetic with (a) supervised finetuning and (b) pretraining only. We experimented with varying values of $\lambda$ (controlling the deviation from the pretrained model, Eq. 3), in particular, between 0.10 and 1.00 with a step size of 0.05. We experimented with fractions of the finetuning dataset (composed of 1024 molecules) ranging from 1% to 100%, with increments of 10%. For task arithmetic, 10,240 negative molecules were used. We evaluated the models by the number of successful design clusters, i.e., the number of design clusters that possess the desired property. This metric combines the evaluation of diversity and accuracy; the higher, the better. The highest number of successful design clusters obtained across $\lambda$ values (task arithmetic), or training subsets (supervised finetuning) was reported for each property. In general, models trained on randomized SMILES yielded more successful design clusters than their canonical counterparts (Fig. A2). Due to the better performance of randomized SMILES, in what follows, we report only the results for randomized SMILES for the zero-shot experiments, while results for canonical SMILES strings were included in the Appendix.
Across design tasks, molecular task arithmetic yielded a higher number of successful design clusters than pretraining and finetuning (9.5K and 11.7K more on average, respectively; Fig. 2a and Fig. A3a ). Except for the $\uparrow$ No. Rings’ task with randomized SMILES, the performance gain by molecular task arithmetic over supervised finetuning is statistically significant across datasets (Mann-Whitney U test, p-value $<$ 0.01). This is a striking finding: despite using no positively labeled molecules, molecular task arithmetic can design a higher number of successful diverse molecules than models fine-tuned on up to 1024 positive compounds. This demonstrates the potential of task arithmetic for generative drug discovery without any positively labeled molecular examples.
To further shed light on the success of molecular task arithmetic, we inspect the success rate (ratio of designs satisfying the task; Fig. 2b and Fig. A3b) and number of total clusters (Fig. 2c and Fig. A3c). Results show that molecular task arithmetic can create, on average, 24.4K more diverse molecule designs than the fine-tuned models and 7.4% more accurate designs than pretraining, while finetuning outperforms molecular task arithmetic by 14.1% in success rate. This is, on one hand, expected, since finetuning has explicit access to molecules fulfilling the desired property. On the other hand, it is surprising that molecular task arithmetic can still create more desired cluster centers with less successful designs. Taken together, these findings display that molecular task arithmetic can preserve the diversity gained by pretraining during conditioning, while finetuning ‘forgets’. Forgetting and loss of generalization with finetuning are observed in different deep learning contexts50, 51, 52, 53, 54, 55, and here we show that task arithmetic can be a solution for molecule design, to maintain higher diversity while achieving accuracy.
Fig. 3: Side effects of transfer learning strategies. Kolmogorov-Smirnov distance between the molecular properties of the designs on SMILES tasks and the pretraining set is computed for off-target properties. Lower distance to the pretraining set indicates that the conditioning caused less change in non-conditioned molecular properties. Bar heights display the means across five training splits, and error bars display the standard deviations.
Side effects.
Another desirable trait of model conditioning is the possibility of having minimal side effects on the ‘off-target’ properties. Being able to minimize side effects would allow for a more precise steering of the model in the chemical space, by preserving desirable molecular properties (i.e., synthesizability, lack of toxicity, etc.). We quantify the side effects of both finetuning and molecular task arithmetic for all design tasks, on the remaining four descriptors. These descriptors have a low correlation with the ‘target’ properties (Table A3). For each task, we computed the Kolmogorov-Smirnov (KS) distance between the off-target molecular properties of the designs and the pretraining set (the lower the KS, the more similar the distributions). The usefulness of this analysis is corroborated by observing the KS distances of the designs obtained by pretraining (Fig. A4), which are all below 5%.
Across the ten design tasks for randomized SMILES, molecular task arithmetic showed a smaller mean distributional distance to the pretraining set in 32 of the 40 comparisons (Fig. 3), of which 15 were confirmed by a statistical test (Mann-Whitney U test, p-value $<$ 0.01). On canonical SMILES strings, all models have a higher average distance (Fig. A5), possibly because canonicalization restricts the models to explore ‘side-effect-free’ portions of the chemical space. Overall, the analysis shows that molecular task arithmetic can combine its accuracy and diversity on the design task, with less side-effect than the traditional finetuning approach. This suggests that task arithmetic allows for a more controlled steering of the models in the desirable portions of the chemical space.
Dual-objective design.
We next explore the capabilities of molecular task arithmetic for dual-objective de novo design. This objective is central in drug discovery, for instance, for multi-target drug design, or selectivity optimization. We define three dual-objective tasks via the molecular properties above: (a) high fraction of sp3-hybridized carbons and high number of hydrogen bond donors, (b) high lipophilicity and low TPSA, and (c) high number of rings and low TPSA. These pairs of properties were selected to be minimally correlated (absolute correlation lower than 0.3; Table A3), so that the models neither face a conflicting nor a trivial task.
To apply molecular task arithmetic in a dual-objective setting (Fig. 1d), we sum the scaled vectors yielding the best performance for each single task (Eq. 3) and scale the final vector in increasing $\lambda$ values (0.05 to 1.00 with a step size of 0.05). For finetuning, we finetuned the models on the training set of the single tasks sequentially (using positive molecules), as in recent literature 56, 57, with both task permutations. We reported the number of (successful) design clusters, success rate, and validity for the $\lambda$ and permutation yielding the highest number of successful design clusters for molecular task arithmetic and finetuning, respectively.
Fig. 4: Zero-shot dual objective molecule design. Models are trained on randomized SMILES representation with sequential supervised finetuning and molecular task arithmetic to design molecules that possess two task properties simultaneously. (a) Number of successful design clusters, (b) number of clusters, (c) success rate, and (d) validity are measured. Experiments are repeated five times, and the mean (bar height) and standard deviation (error bar) are reported.
[…]
4.3 Effect of Task Vector Scaling
Here, we further dig into molecular task arithmetic, to study the impact of $\lambda$ (Eq. 3) on the performance. The parameter $\lambda$ scales the task vector length and determines how far the task model ‘lands’ from the pretrained model after task negation. We compute validity and success rate across all 20 design tasks with 19 increasing $\lambda$ values (from $0.05$ to $1.00$ with a step of $0.05$). We measured the validity and success rate in those increasing lengths of task vectors, to reveal how much the model can move in the weight space before its generation capabilities collapse.
Across experiments, the validity decreases with increasing $\lambda$ values (Fig. 5 and Fig. A7). We explain this by the objective of the pretraining task: since maximizing next token prediction accuracy optimizes weights for generation validity, moving the model away causes a drop. The success rate also displays consistent patterns (Fig. 5 and Fig. A7): it first increases, then plateaus (typically for $0.30\leq\lambda\leq 0.50$), and monotonically decreases afterward. This shows that molecular task arithmetic gradually conditions the pretrained model as the model moves along the task direction, and then, generative capabilities start limiting the number of successful designs. The consistent behaviors offer a practical insight to tune $\lambda$ for molecular task arithmetic models. Models can be edited by setting $\lambda=0.30$ and gradually increasing the value until the success rate starts decreasing. Thanks to the peaking behavior, this heuristic can find the $\lambda$ value optimal for the task.
Fig. 5: Impact of molecular task vector scaling. For each design task defined on randomized SMILES strings, molecular task arithmetic is applied in 19 increasing scaling factors ($\lambda$; Eq. 3). 100,000 molecules are designed with all values of $\lambda$ and validity and success rate are measured across five training splits. Lines denote the means across runs, and shaded areas describe the standard deviations.
4.4 Boosting Few-Shot Molecule Design
Fig. 6: Few-shot molecule design with molecular task arithmetic. Models are trained on an increasing number of positive data points with supervised finetuning and molecular task arithmetic. Molecular task arithmetic is first applied using negative data, and then the resulting model is fine-tuned with the available task data. The mean and the standard deviation of the number of successful design clusters across five splits are computed (lines and shaded regions, respectively).
We next delve into another interesting problem in drug discovery: few-shot molecule design. In few-shot molecule design, some positive labeled molecules are already available, which are used for molecule design. Additionally, negative molecules are usually available (and in general more abundant), but are not leveraged in traditional finetuning pipelines. Here we set out to investigate whether molecular task arithmetic can allow leveraging negative molecules alongside positive ones to boost the capacity of models to generate molecules with desirable properties.
We applied molecular task arithmetic as in the zero-shot design experiments (Eq. 3) first, and then fine-tunex the obtained model with the known positive molecules (Fig. 1e). In other words, this pipeline replaces the pretraining model used for supervised finetuning with the molecular task arithmetic model, to harness the benefits of abundant and structurally diverse negative molecules. We compare the proposed approach to traditional supervised finetuning by computing the number of design clusters when increasing the number of positive molecules.
Across all 20 tasks and the number of positive molecules studied, molecular task arithmetic surpassed the number of successful clusters obtained with finetuning, up to a 10K increase (Fig. 6). A deeper look reveals that molecular task arithmetic can design at least as many successful designs as supervised finetuning (Fig. A8), and yet maintain the design diversity (Fig. A9). Overall, these experiments point to a synergy: whenever positive molecules are available, molecular task arithmetic can be combined with supervised finetuning to increase the number of diverse hits.
Molecular task arithmetic can reach maximum performance even with fewer than 50 positive molecules. The number of clusters gradually decreases as the model is finetuned with further molecules. Such performance fluctuations, however, have a smaller magnitude than vanilla finetuning, underlining the robustness of molecular task arithmetic. Taken together, the experiments highlight molecular task arithmetic as a data-efficient, robust, and well-performing transfer learning strategy for few-shot molecule design. Combined with its simplicity, molecular task arithmetic can potentially become the new standard for transfer learning in de novo design, by allowing one to leverage both positive and negative data simultaneously.
5 Related Work
Task arithmetic.
Discovering semantic directions in the model weight space has emerged as a new research direction that offers opportunities to edit model behavior with simple arithmetic operations 25, 27, 28, 29. Named as task arithmetic, this perspective has enabled higher performance with model merging 26, in-context task learning28, and zero-shot information retrieval27. Its capability to recognize task directions from finetuning of opposite labels offers potential for molecular field, where negative samples are dominantly more available than positive ones. Here we explored this potential.
Molecule design.
Deep learning for de novo drug design has been an active field in the last decade 42, 58, 59. Early approaches trained sequence models on string representations of molecules with a next-token prediction task, using language modeling42, 58, 59. Recent works used graph and 3D-based representations to represent molecules and graph neural networks and diffusion for learning 60, 61, 62, 63, 64, 65. While these approaches enable more fine-grained representation and high steerability in learning, decoder-only language models of molecular strings, termed chemical language models, remain popular to date due to their simplicity and performance40. Furthermore, one molecule can have multiple SMILES strings, which enables data augmentation for deep learning in low data 34, 56, 66, 45.
6 Limitations
Molecular task arithmetic relies on negative data. While negative molecules are abundant due to the characteristics of the chemical space, this still forms a limitation to applying molecular task arithmetic. Our experiments also revealed that the validity of generations drops with molecular task arithmetic. Smarter weight edits29 and/or alternative architectures should be studied to minimize this side effect. Last, this study tested molecular task arithmetic only for de novo design using LSTMs on SMILES strings. Experiments on other molecular generation tasks with more architectures and representations are needed to determine the limits of molecular task arithmetic as a new transfer learning strategy for deep learning in chemistry.
In this study, molecular task arithmetic was tested on molecular properties that can be exactly calculated and that vary linearly. It is yet to be determined whether negating more complex and non necessarily linear properties (e.g., inactivity, in the case of ligand-protein interaction) would yield molecules with the desired properties (e.g., biological activity).
7 Conclusion
Drug discovery datasets are usually small and lack structural diversity. While deep generative modeling and transfer learning have already been transformative for drug design, they still must face this inherent challenge. This work proposed a remedy: molecular task arithmetic. Molecular task arithmetic is a new transfer learning strategy that leverages the diverse and abundant negative molecules. Our approach unlocks zero-shot ligand-based molecule design in single- and dual-objective settings – creating a new paradigm for drug discovery. Experiments across ten design tasks show that molecular task arithmetic – which uses no positive examples – can design more diverse hits that fulfill desirable properties than models trained solely on positive data. We show that molecular task arithmetic can also leverage the available positive molecules to further propel chemical space exploration – outperforming traditional transfer learning across all design tasks. Thanks to its simplicity and performance, molecular task arithmetic bears the potential to become a well-established transfer learning strategy for drug discovery, a field where negative data abound.
CRediT Author Contributions
Conceptualization: RÖ and FG; Data curation: RÖ and SdR; Formal analysis: RÖ and SdR; Funding acquisition: FG; Investigation: all authors; Methodology: all authors; Visualization: RÖ, with contributions from all authors; Writing – original draft: RÖ; Writing – review & editing: RÖ and FG.
Acknowledgments and Disclosure of Funding
This research was co-funded by the European Union (ERC, ReMINDER, 101077879, to F.G.). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This publication is also part of the project ChemWISE with file number VI.Vidi.233.164 of the research programme Vidi ENW, which is (partly) financed by the Dutch Research Council (NWO) under the grant: https://doi.org/10.61686/IVDFS18985 (to F.G.). The authors are grateful to the Centre for Living Technologies for financial support, and to the Netherlands Organization for Scientific Research for compute resources (grant EINF-7609 to R.Ö.). The authors thank Elena Frasnetti for her contribution to model pretraining and Selen Parlar Özçelik for her feedback on the manuscript. The authors also acknowledge the Molecular Machine Learning team at TU/e for fruitful scientific discussions.
Appendix
Experimental settings
A chemical language model with an LSTM backbone is pretrained on a previously curated dataset of 1.5M drug-like molecules from ChEMBL[21, 67]. 100 models were trained for either canonical or randomized SMILES strings, with random hyperparameters (Table A1). For each representation, the model that generates the highest number of unique and new designs was selected for the finetuning stage.
The finetuning sets were curated from a previous study [24] by ensuring that no finetuning molecule is present in the pretraining set. The thresholds for high and low values are computed by rounding the median of the molecular property values among the pretraining molecules (Table A2 and Fig. A1). Five molecule sets are curated for each task (1024 train, 256 validation, 256 test molecules per split), where each molecule carries the task property. For models trained on randomized SMILES, 10-fold SMILES augmentation is used during finetuning. Early stopping on validation loss (cross-entropy) with a patience of five epochs and loss tolerance of $1\times 10^{-5}$ is employed. 100,000 molecules were generated with each model.
Four metrics were computed to evaluate the quality of the de novo designs across the study:
  • Validity: the percentage of ‘chemically valid’ molecules across designs. Chemical validity is checked by attempting to create a molecule object in rdkit, after filtering out empty string designs.
  • Success rate: the frequency of distinct molecules outside the training sets that possess the task property. Molecules are deduplicated via canonicalizing the designs and keeping only one of the identical strings. Success rate evaluates the accuracy of the model (the higher, the better).
  • Number of clusters: the number of clusters identified by the sphere exclusion algorithm. Tanimoto distance on extended connectivity fingerprints [68] is used with a threshold of 65% to identify the number of distinct clusters. The metric is linked to #Circles [69], and the rdkit implementation is used. The higher the number of clusters, the more diverse the design library – a desired trait for drug discovery [70].
  • Number of successful molecule clusters: the number of cluster centers that possess the task property. This metric combines the evaluation of diversity and accuracy, and is used as the main evaluation metric across the study. Higher values are preferred.
Table A1: Hyperparameter space for pretraining the LSTMs. 100 models are trained by randomly subsampling the hyperparameter space. NVIDIA A100 GPUs with 40GB of memory are used. Training one model on 1.5M SMILES strings took approximately 12 hours on average. Models with different parameters are trained simultaneously by using one GPU per model in a supercomputer.
Hyperparameter name Space
Number of layers 1, 2, 4, 6, 8
Hidden state dimension 256, 512, 1024, 2048
Dropout rate 0.0, 0.1, 0.15, 0.2, 0.25
Vocabulary size 33
Input sequence length 82
Learning rate $1\times 10^{-4}$, $5\times 10^{-4}$, $1\times 10^{-3}$, $5\times 10^{-3}$, $1\times 10^{-2}$
Maximum number of epochs 1000
Batch size 8192
Optimizer Adam
Table A2: Thresholds used to define the design tasks. Values less than or equal to the threshold are considered low, and molecules with higher values are labeled high.
Property Threshold
Fraction of sp3-hybridized carbons 0.3
Number of hydrogen bond donors 1
Number of rings 3
Octanol-water partition coefficient (logP) 3.5
Topological polar surface area (TPSA) 75
Fig. A1: Distribution of molecular properties in the pretraining set. The dashed line corresponds to the median, which is also annotated.
Fig. A2: Comparison of SMILES representations. The average number of successful design clusters is computed per design task using canonical and randomized SMILES representations. Bar heights report the scores obtained by using canonical SMILES subtracted from using randomized SMILES.
Fig. A3: Zero-shot molecule design with molecular task arithmetic. Models were trained on 10 design tasks of canonical SMILES strings, with molecular task arithmetic and supervised finetuning. 100,000 molecules were designed and clustered. (a) The number of cluster centers that possess the desired properties, (b) ratio of the designs that satisfies the design task, and (c) number of clusters were computed. The pretrained model is also included in the analysis as a baseline. Bar heights report the mean statistics across five training sets, and error bars denote the standard deviation. Finetuning was conducted up to the training sets of 1024 molecules, while molecular task arithmetic used 10,240 negative molecules and no labeled positive data. Yet, molecular task arithmetic obtained higher number of successful design clusters across design tasks.
Fig. A4: Side effects of conditioning strategies and pretraining. Compuations are described in (Fig. 3). Bar heights display the means across five training splits, and error bars display the standard deviations. In 32 of 40 comparisons, molecular task arithmetic caused less side-effects.
Fig. A5: Side effects of transfer learning strategies on canonical SMILES experiments. Kolmogorov-Smirnov distance between the molecular properties of the designs on canonical SMILES tasks and the pretraining set is computed, except for the task property. Bar heights display the means across five training splits, and error bars display the standard deviations.
Table A3: The correlations of molecular properties. Pearson correlations are computed over the pretraining set. Correlations for the tasks in dual-objective design experiments are marked with boldface.
Property No. H. Donor No. Rings logP TPSA Frac. sp3 C
No. H. Donor - -0.09 -0.28 0.17 0.02
No. Rings -0.09 - 0.40 0.26 0.13
logP -0.28 0.40 - -0.12 0.08
TPSA 0.17 0.26 -0.12 - 0.57
Frac. sp3 C 0.02 0.13 0.08 0.57 -
Fig. A6: Zero-shot dual objective molecule design on canonical SMILES. Models are trained on canonical SMILES representation with sequential supervised finetuning and molecular task arithmetic to design molecules that possess two task properties simultaneously. Number of successful design clusters, success rates, number of clusters, and validity are measured. Experiments are repeated five times, and the mean (bar height) and standard deviation (error bar) are reported.
Fig. A7: Impact of molecular task vector scaling on canonical SMILES. For each design task defined on canonical SMILES strings, molecular task arithmetic is applied in increasing scaling factors ($\lambda$; Eq. 2). 100,000 molecules are designed with all $\lambda$s and validity and success rate are measured across five training splits. Lines denote the means across runs, and shaded areas describe the standard deviations.
Fig. A8: Success rates for few-shot molecule design. Experimental pipeline is detailed in (Fig. 6). The mean and the standard deviation of the number of successful design clusters across five splits are computed (lines and shaded regions, respectively).
Fig. A9: Number of design clusters for few-shot molecule design. Experimental pipeline is detailed in (Fig. 6). The mean and the standard deviation of the number of successful design clusters across five splits are computed (lines and shaded regions, respectively).